OpenStudy (anonymous):

When the quadratic equation y = (x − 2)(3x + 1) is written in standard form, what is the value of the coefficient "b"? −5 3 −2 5

OpenStudy (anonymous):

-5 is the correct answer. y=(x-2)(3x+1) y=3x^2+x-6x-2 y=3x^2-5-2 therefore b=-5

OpenStudy (anonymous):

thank you!